OpenStudy (anonymous):

Can anyone help me with this summation problem: 5+7+9+11+13=sigma top number is 4 the index starts at 0. I'm supposed to find a function, but don't know how to solve that. Thank you.

OpenStudy (anonymous):

look like you are adding two each time

OpenStudy (anonymous):

\(4\times 2+5=13\) so perhaps try \[\sum_0^42k+5\]

OpenStudy (anonymous):

\[\sum_{i=0}^{4}5+2i=5+7+9+11+13\]The reason it's 5 + 2i is because 5+2(0) = 5, 5+2(1)=7, and so on.

OpenStudy (anonymous):

or \(i\), either way

OpenStudy (anonymous):

I see. So in order to get the equation one must look at the patter at the numbers being added then multiply by the n of the summation and add it to the first number to get the last one?